OpenStudy (bimbim2697):

A(n)=A(1)+(n-1)d 1)YOU HAVE A GIFT CARD FOR A COFFE SHOP WORTH $90.EACH DAY YOU USE THE CARD TO GET COFFEE FOR $4.10.WRITE A RULE TI REPRESENT THE AMMT.OF MONEY LEDT ON THE CARD AS AN ARITHMETIC SEQUENCE .WHAT IS THE VALUE OF THE CARD AFTER BUYING 8 COFFEES

myininaya (myininaya):

So if you have $90 and it is the first day you use your card, you have wait left 90-4.10. Second day you have 90-4.10-4.10=90-2(4.10) Third day you have 90-4.10-4.10-4.10=90-3(4.10) ... so on the nth day what do you have on your card?

myininaya (myininaya):

what is the pattern?

OpenStudy (bimbim2697):

-4.10 each day

myininaya (myininaya):

1st day we have 90-1(4.10) 2nd day we have 90-2(4.10) 3rd day we have 90-3(4.10) 4th day we have 90-4(4.10) ... nth day we have ?

OpenStudy (bimbim2697):

90-9(4.10)

OpenStudy (bimbim2697):

So this would be the rule right?

myininaya (myininaya):

90-n(4.10) is what we have on the nth day

OpenStudy (bimbim2697):

$53.10

myininaya (myininaya):

now if you want to figure out what you have left on that gift card on the 8 th day you just plug in your little 8 in the rule i gave you

OpenStudy (bimbim2697):

so the rule in 90-n(4.10)

myininaya (myininaya):

yes you gave me the 9 th day that is the nth day i was asking for

OpenStudy (bimbim2697):

oh thank u so much can u help me with the other one please